CH4 + 2O2 = CO2+2H2O

If 312g of CO2 is produced what mass of CH4 is required?


1
Expert's answer
2021-02-28T06:20:02-0500

"CH_4+2O_2 \rightarrow CO_2+2H_2O"


Weight of "CO_2" given "=" 312 gm

moles of "CO_2" "(n_{CO_2})=" "\dfrac{weight of CO_2}{Molar mass of CO_2}"


"n_{CO_2}=\dfrac{312}{44}=7.09" moles "=7.10" moles (approx.)


As from the stoichiometry we can easily see that 1 mol of "CO_2" achieved from 1 mol of "CH_4" ,so "7.10" moles achieved from "7.10" moles of "CH_4"


So moles of "CH_4=7.10" moles

mass of "CH_4=" moles of "CH_4\times" molar mass of "CH_4"

"=7.10\times16=113.16" gm